### v2

# On the Distribution of Points of Valuation $1$ for a Polynomial in Two Variables

Abstract
We investigate the variation in the total number of points in a random $p\times p$ square in $\mathbb{Z}^{2}$ where the $p$-adic valuation of a given polynomial in two variables is precisely $1$. We establish that this quantity follows a Poisson distribution as $p\rightarrow\infty$ under a certain conjecture. We also relate this conjecture to certain uniform distribution properties of a vector valued sequence.
1 Introduction
The $p$-adic valuation of integer sequences is a fundamental concept in number theory, providing insight into the divisibility properties of integers with respect to a prime number $p$. The $p$-adic valuation $\nu_{p}(n)$ of an integer $n$ is defined as the largest exponent $e$ such that $p^{e}$ divides $n$. Historically, the concept of $p$-adic valuation stems from the work of Kurt Hensel[6], who introduced $p$-adic numbers in 1897. This development allowed mathematicians to study numbers and their divisibility in a local sense, focusing on individual primes. Over time, the theory of $p$-adic numbers and valuations was expanded by Emil Artin, Alexander Grothendieck, and others, making it a crucial tool in areas such as Diophantine equations and algebraic geometry.
One essential result in $p$-adic valuation is Legendre’s formula[7], which calculates the $p$-adic valuation of $n!$. For a prime $p$, the formula is given by:
$\nu_{p}(n!)=\sum_{k=1}^{\infty}\left\lfloor\frac{n}{p^{k}}\right\rfloor\cdot$
This sums the number of multiples of $p,p^{2},p^{3},\dots$ less than or equal to $n$, capturing how many times $p$ divides $n!$. This formula has important applications in combinatorics, particularly in understanding binomial coefficients and their divisibility by primes.
The work presented in [3] forms part of a general project initiated by Moll and his co-authors to analyse the set
$V_{x}=\{\nu_{p}(x_{n}):n\in\mathbb{N}\}$,
for a given sequence $x=\{x_{n}\}$ of integers. The $2$-adic valuation of $\{n^{2}-a\}$ is studied in [3]. It is shown that $n^{2}-a,a\in\mathbb{Z}$ has a simple closed form[2, 4] when $a\not\equiv 4,7\mod 8$. For these two remaining cases the valuation is quite complicated. It is studied by the notion of the valuation tree. In work by the second author [9], the set
$V_{f}$ = $\{\nu_{p}(f(m,n)):m,n\in\mathbb{N}\}$,
where $f(x,y)\in\mathbb{Z}[x,y]$, is discussed by the generalized notion of the valuation tree.
For a fixed polynomial $f(x,y)\in\mathbb{Z}[x,y]$, let $C_{p}=\{(x,y)\in\mathbb{Z}^{2}:\nu_{p}(f(x,y))>0\textit{ and }0\leq x,y<p\}$ be the points of the curve given by the equation $f(x,y)=0$ over $\mathbb{F}_{p}$ and $m=|C_{p}|$ be the number of points in $C_{p}$.
In order to study the points in $\mathbb{Z}^{2}$ where the valuation of $f$ is exactly $1$, we observe the patterns occurring for various polynomials $f$ as shown in Figure 1.
Figure 1: Scatter Plot of $(x,y)$ satisfying: Upper Left: $\nu_{p}(x^{3}+y^{3}+x^{2}y+y+1)=1$, for $p=17$; Upper Right: $\nu_{p}(x^{3}+y^{2}x+xy+x+y+1)=1$, for $p=19$; Lower Left: $\nu_{p}(x^{3}+xy+x+y+1)=1$, for $p=23$; Lower Right: $\nu_{p}(y^{2}x+xy+x+y+1)=1$, for $p=29$.
These seem to encode some elements of randomness which we try to quantify further by studying the properties of indicator random variables $Y_{x,y}$ defined in Section 3. To summarize, these variables are identically distributed Bernoulli random variables with probability $1/p$ and are pairwise independent most of the time.
In a different direction, we study the variation in total number of points with fixed valuation in a random square. More precisely, define the random variable $X:\mathbb{F}_{p}\times\mathbb{F}_{p}\rightarrow\mathbb{Z}$ where $X(k,l)$ is the total number of points $(x,y)\in\{kp,kp+1,\ldots,kp+p-1\}\times\{lp,lp+1,\ldots,lp+p-1\}$ such that $\nu_{p}(f(x,y))>1$. In other words, $X$ represents the number of points in the random $p\times p$ square where the valuation of $f$ is greater than $1$. Observing the distribution of $X$ for various primes $p$, we get the following Figure 2.
Figure 2: Comparison of the Poisson distribution ($\lambda=1$) with the distribution of $X$ for $f(x,y)=x^{3}+y^{2}+xy+1$ and the prime numbers, Upper Left: $p=211$; Upper Right: $p=503$; Lower Left: $p=1013$; Lower Right: $p=1033$.
This behaviour for large $p$ can be explained by the following reasoning under the assumption that the $Y_{x,y}$ are mutually independent: By (21), we get that $X$ follows the Binomial distribution with parameters $m$ and $1/p$. Note that their product $m/p\rightarrow 1$ as $p\rightarrow\infty$ by the Hasse-Weil bounds [10] and hence the distribution of $X$ tends to a Poisson distribution with parameter $1$.
Now we are in position to state our main result.
Theorem 1.
Assume that Conjecture 1 is satisfied. Then X converges in distribution to a Poisson random variable with parameter $\lambda=1$ as $p\rightarrow\infty$.
Before stating Conjecture 1, we would need some notations. For any $k$-tuple $\mathbf{P}=((r_{1},s_{1}),(r_{2},s_{2}),\dots,(r_{k},s_{k}))$ of points on the curve $C_{p}$, denote the associated matrix $B_{k}(\mathbf{P})$ to be
$B_{k}(\mathbf{P})=\begin{bmatrix}f_{x}(r_{1},s_{1})&f_{y}(r_{1},s_{1})&f(r_{1},s_{1})/p\\ f_{x}(r_{2},s_{2})&f_{y}(r_{2},s_{2})&f(r_{2},s_{2})/p\\ \vdots&\vdots&\vdots\\ f_{x}(r_{k},s_{k})&f_{y}(r_{k},s_{k})&f(r_{k},s_{k})/p\end{bmatrix}.$ (1)
Let $m_{k,1}$ denote the number of element-wise distinct $k$-tuples of points on $C_{p}$ whose associated matrix $B_{k}(\mathbf{P})$ has rank $1$. Let $m_{k,2}$ denote the number of element-wise distinct $k$-tuples for which $B_{k}(\mathbf{P})$ has rank $2$ and the first $2$ columns of $B_{k}(\mathbf{P})$ are linearly independent.
For $k=1$, we note that $m_{k,2}=0$ trivially and $m_{k,1}=m$, the number of points on $C_{p}$. For $k\geq 2$, we note (see the proof of Corollary 2) that $m_{k,1}$ is bounded by a constant depending only on $f$. Regarding $m_{k,2}$, we have the following conjecture.
Conjecture 1.
For all $k\geq 2$, we have
$\lim_{p\to\infty}m_{k,2}/p^{2}=1.$ (2)
Numerical observations appear to corroborate this conjecture as seen in Figure 3.
Figure 3: Left: Plot of $m_{2,2}/p^{2}$ versus $p$; Right: Plot of $m_{3,2}/p^{2}$ versus $p$ for the polynomial $f(x,y)=x^{3}+y^{2}+xy+1$.
We also show in Section 4 that this conjecture is a natural consequence of certain versions of uniform distribution of $\mathcal{F}_{p}$ in $\mathbb{F}_{p}^{3}$, where
$\mathcal{F}_{p}=\{(f_{x}(r,s),f_{y}(r,s),f(r,s)/p)\in\mathbb{F}_{p}^{3}\ |\ (r,s)\in C_{p}\}.$
We now describe the organization of this paper. In Section 2, we prove several elementary results on the number of solutions to certain valuation questions. In Section 3, we interpret these results as describing the single and joint distribution of indicator random variables indexed by points in $C_{p}$. We also derive the implication of Conjecture 1 on the sum of expectation of $k$-fold products of these random variables. Section 4 contains results on equidistribution of $\mathcal{F}_{p}$ as mentioned above. Section 5 contains the proof of Theorem 1.
2 Preliminaries
In this section, we study about the cardinality of the set $\{(x,y):\nu_{p}(f(x,y))=1,0\leq x,y<p^{2}\}$ and find the following results.
Proposition 1.
Let $f(x,y)\in\mathbb{Z}[x,y]$ be an integer polynomial with non-zero constant term and $(x,y)\in C_{p}$ such that $f_{x}(x,y)\not\equiv 0\pmod{p}$. Then the cardinality of the set $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(x+kp,y+lp))=1\}$ is $p(p-1)$.
Proof.
Let $(x,y)\in\{0,1,...,p-1\}^{2}$ such that $f(x,y)\equiv 0\pmod{p}$ and $f(x,y)\not\equiv 0\pmod{p^{2}}$. Now observe that any point $x\in\{0,1,...,p^{2}-1\}$ is in the set $\{i+jp:0\leq i,j\leq p-1\}$. Consider
$\begin{split}&f(x+kp,y+lp)\equiv 0\pmod{p^{2}}\end{split}$ (3)
where $(k,l)\in\{0,1,2,...,p-1\}$. Now by Taylor’s series expansion, we have
$f(x,y)+kpf_{x}(x,y)+lpf_{y}(x,y)\equiv 0\pmod{p^{2}}$ (4)
We know that $f(x,y)\equiv 0\pmod{p}$ which implies there exist an $\alpha$ such that $f(x,y)=\alpha p$. Now by using (4), we have
$\alpha+kf_{x}(x,y)+lf_{y}(x,y)\equiv 0\pmod{p}.$ (5)
Then we have
$k\equiv\frac{-\alpha-lf_{y}(x,y)}{f_{x}(x,y)}\pmod{p}.$ (6)
So $(k,l)$ has $p$ choices such that $f(x+kp,y+lp)\equiv 0\pmod{p^{2}}$. Hence $(k,l)$ has $p^{2}-p$ choices such that $f(x+kp,y+lp)\not\equiv 0\pmod{p^{2}}$.
Corollary 1.
Let $m$ and $s$ be the cardinalities of the sets $C_{p}$ and $S$ respectively, where $S=\{(x,y):\nu_{p}(f(x,y))=1,0\leq x,y<p^{2}\}$. Then $s=mp(p-1)$.
We examine the mutual influence of points $(r_{1},s_{1})$ and $(r_{2},s_{2})$ on the curve $C_{p}$. Assume that
$A=\begin{pmatrix}f_{x}(r_{1},s_{1})&f_{y}(r_{1},s_{1})&f(r_{1},s_{1})/p\\ f_{x}(r_{2},s_{2})&f_{y}(r_{2},s_{2})&f(r_{2},s_{2})/p\end{pmatrix}$ (7)
and $r$ is the rank of $A$. In the case that $r=2$ and the first 2 columns of $A$ are linearly independent, we establish the following
Proposition 2.
Let $f(x,y)\in\mathbb{Z}[x,y]$ be an integer polynomial with non-zero constant term and $f(r_{1},s_{1})\equiv 0\pmod{p}$ and $f(r_{2},s_{2})\equiv 0\pmod{p}$. Suppose that the first two columns of $A$ are linearly independent. Then the cardinality of the set
  • 1.
    $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))=1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))=1\}$ is $(p-1)^{2}$.
  • 2
    $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))>1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))=1\}$ is $p-1$.
  • 3
    $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))>1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))>1\}$ is $1$.
Proof.
Let $(r_{1},s_{1}),(r_{2},s_{2})\in C_{p}$ such that $\nu_{p}(f(r_{1},s_{1}))=1$ and $\nu_{p}(f(r_{2},s_{2}))=1.$ Now observe that any point $x\in\{0,1,...,p^{2}-1\}$ is in the set $\{i+jp:0\leq i,j\leq p-1\}$. For $i,j\in\{0,1,2,...,p-1\}$, consider the following Taylor’s series expansion,
$f(r_{1}+ip,s_{1}+jp)\equiv f(r_{1},s_{1})+ipf_{x}(r_{1},s_{1})+jpf_{y}(r_{1},s_{1})\pmod{p^{2}}$ (8)
We know that $f(r_{1},s_{1})\equiv 0\pmod{p}$ which implies there exist $\alpha,\alpha^{\prime}$ such that $f(r_{1},s_{1})=\alpha p$ and $f(r_{1}+ip,s_{1}+jp)=\alpha^{\prime}p$. So by using equation (8), we have
$if_{x}(r_{1},s_{1})+jf_{y}(r_{1},s_{1})\equiv-\alpha+\alpha^{\prime}\pmod{p}.$ (9)
Similarly by using $\nu_{p}(f(r_{2},s_{2}))=1$ we have
$if_{x}(r_{2},s_{2})+jf_{y}(r_{2},s_{2})\equiv-\beta+\beta^{\prime}\pmod{p},$ (10)
for integers $\beta,\beta^{\prime}$ such that $f(r_{2},s_{2})=\beta p$ and $f(r_{2}+ip,s_{2}+jp)=\beta^{\prime}p$. By using equations (9) and (10) we get the following system of two equations in variables $i,j$
$A\begin{pmatrix}i\\ j\\ 1\end{pmatrix}\equiv\begin{pmatrix}\alpha^{\prime}\\ \beta^{\prime}\end{pmatrix}\pmod{p}.$ (11)
Now we know that the $f_{x}(r_{1},s_{1})f_{y}(r_{2},s_{2})-f_{y}(r_{1},s_{1})f_{x}(r_{2},s_{2})\not\equiv 0\pmod{p}$ so the above system of equation has a unique solution for each $\alpha^{\prime}$ and $\beta^{\prime}$.
Now the set $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))=1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))=1\}$ corresponds to $\alpha^{\prime}\not\equiv 0\pmod{p}$ and $\beta^{\prime}\not\equiv 0\pmod{p}$. Hence its cardinality is $(p-1)^{2}$. Similarly the set $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))>1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))=1\}$ corresponds to $\alpha^{\prime}\equiv 0\pmod{p}$ and $\beta^{\prime}\not\equiv 0\pmod{p}$. So this set has $p-1$ elements. Finally, the set $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))>1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))>1\}$ corresponds to $\alpha^{\prime}\equiv 0\pmod{p}$ and $\beta^{\prime}\equiv 0\pmod{p}$. So it has only one element. ∎
Next, in the case that the first $2$ columns of $A$ are linearly dependent, we establish the following.
Proposition 3.
Let $f(x,y)\in\mathbb{Z}[x,y]$ be an integer polynomial with non-zero constant term and $f(r_{1},s_{1})\equiv 0\pmod{p}$ and $f(r_{2},s_{2})\equiv 0\pmod{p}$. Suppose that first two columns of $A$ are linearly dependent. Then the cardinality of the set
  • 1.
    $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))=1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))=1\}$ is $p(p-1)$ if $r=1$ and $p(p-2)$ if $r=2$.
  • 2.
    $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))>1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))=1\}$ is 00 if $r=1$ and $p$ if $r=2$.
  • 3.
    $\{(k,l)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+kp,s_{1}+lp))>1\textit{ and }\nu_{p}(f(r_{2}+kp,s_{2}+lp))>1\}$ is $p$ if $r=1$ and 00 if $r=2$.
Proof.
We begin the proof by noting that equation (11) holds in this case too, as all the steps in the proof of Prop. 2 leading up to (11) are valid. We also assume that the second column of $A$ is a multiple of the first column by $a\pmod{p}$ as the first $2$ columns are given to be linearly dependent.
  • Case 1.
    The set $\{(i,j)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+ip,s_{1}+jp))=1\textit{ and }\nu_{p}(f(r_{2}+ip,s_{2}+jp))=1\}$ corresponds to $\alpha^{\prime}\not\equiv 0\pmod{p}$ and $\beta^{\prime}\not\equiv 0\pmod{p}$. If rank of $A$ is $1$ then $\alpha\equiv a\beta\pmod{p}$ and the system (11) has solution if $\alpha^{\prime}\equiv a\beta^{\prime}\pmod{p}$. So the system (11) has $p$ solutions for each fixed $\alpha^{\prime}$. But $\alpha^{\prime}$ has $p-1$ choices. Hence in this case, the system (11) has total $p(p-1)$ solutions.
    In the other case when rank of $A$ is $2$, then $\alpha\not\equiv a\beta\pmod{p}$. So the system (11) has solution if $\alpha^{\prime}\not\equiv a\beta^{\prime}\pmod{p}$. Hence $\alpha^{\prime}$ has $p-2$ choices. So the system (11) has total $p(p-2)$ solutions.
  • Case 2.
    The set $\{(i,j)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+ip,s_{1}+jp))>1\textit{ and }\nu_{p}(f(r_{2}+ip,s_{2}+jp))=1\}$ corresponds to $\alpha^{\prime}\equiv 0\pmod{p}$ and $\beta^{\prime}\not\equiv 0\pmod{p}$. So system (11) becomes
    $A\begin{pmatrix}i\\ j\\ 1\end{pmatrix}\equiv\begin{pmatrix}0\\ \beta^{\prime}\end{pmatrix}\pmod{p}.$ (12)
    Hence the system 12 has 00 solutions if rank of $A$ is $1$ and $p$ solutions when rank of $A$ is $2$.
  • Case 3.
    The set $\{(i,j)\in\mathbb{F}_{p}^{2}:\nu_{p}(f(r_{1}+ip,s_{1}+jp))>1\textit{ and }\nu_{p}(f(r_{2}+ip,s_{2}+jp))>1\}$ corresponds to $\alpha^{\prime}\equiv 0\pmod{p}$ and $\beta^{\prime}\equiv 0\pmod{p}$. So system (11) becomes
    $A\begin{pmatrix}i\\ j\\ 1\end{pmatrix}\equiv\begin{pmatrix}0\\ 0\end{pmatrix}\pmod{p}.$ (13)
    Hence the system 13 has $p$ solutions if rank of $A$ is $1$ and 00 solutions if rank of $A$ is $2$.
3 Indicator random variables for valuation greater than $1$
For each $(x,y)\in C_{p}$, define a random variable $Y_{x,y}:\mathbb{F}_{p}\times\mathbb{F}_{p}\rightarrow\{0,1\}$ as follows
$Y_{x,y}(k,l)=\begin{cases}1,&\nu_{p}(f(x+kp,y+lp))>1,\\ 0,&\text{otherwise}.\end{cases}$
In other words, $Y_{x,y}$ is an indicator function on the translates of $(x,y)$ where $f$ has $p$-adic valuation greater than 1. By using Proposition 1, we know that the probability mass function of $Y_{x,y}$ is
$P(Y_{x,y}=i)=\begin{cases}1/p,&i=1,\\ (p-1)/p,&i=0.\end{cases}$
Thus the random variable $Y_{x,y}$ has following properties:
  1. (i)
    The expectation $E[Y_{x,y}]=1/p.$
  2. (ii)
    The variance $Var(Y_{x,y})=(p-1)/p^{2}$.
  3. (iii)
    The $r^{\mathrm{th}}$ moment about origin $\mu_{r}^{\prime}=1/p$.
  4. (iv)
    The $r^{\mathrm{th}}$ central moment $\mu_{r}=\big{(}(p-1)^{r}+(-1)^{r}(p-1)\big{)}/p^{r+1}.$
Further, $Y_{x,y}$ has finite mean and variance, and so it satisfies the Chebyshev’s Inequality
$P(|Y_{x,y}-1/p|\geq\epsilon)\leq\dfrac{p-1}{\epsilon^{2}p^{2}},$
where $\epsilon$ is a positive real number.
For the pair of points of the curve $C_{p}$ satisfying the conditions of Prop. 2, the joint probability mass function is
$P(Y_{x,y}=i,Y_{r,s}=j)=\begin{cases}(p-1)^{2}/p^{2},&i=j=1,\\ (p-1)/p^{2},&i=1,j=0\text{ or }i=0,j=1,\\ 1/p^{2},&i=j=0.\end{cases}$
where $(r,s)\neq(x,y)\in C_{p}$. Clearly $P(Y_{x,y}=i,Y_{r,s}=j)=P(Y_{x,y}=i)P(Y_{r,s}=j)$ for all $i,j$. So these are independent random variables.
Similarly we can define the following joint probability mass functions for the pair of points satisfying the conditions of Prop. 3
$P(Y_{x,y}=i,Y_{r,s}=j)=\begin{cases}(p^{2}-2p)/p^{2},&i=j=1,\\ 1/p,&i=1,j=0\text{ or }i=0,j=1,\\ 0,&i=j=0.\end{cases}$
in the case $r=2$. In the other case $r=1$, we have
$P(Y_{x,y}=i,Y_{r,s}=j)=\begin{cases}(p^{2}-p)/p^{2},&i=j=1,\\ 0,&i=1,j=0\text{ or }i=0,j=1,\\ 1/p,&i=j=0.\end{cases}$
Clearly, these are not independent random variables in either case. However, the correlation coefficient is $-1/(p-1)$ for $r=2$ which is the majority case and hence they are close to being uncorrelated.
For ease of notation, we now fix any order on $C_{p}$ for the rest of this article and denote $Y_{i}$ to be the random variable corresponding to the $i^{\mathrm{th}}$ point.
Recall from Section 1, that $m_{k,2}$ denotes the number of element-wise distinct $k$-tuples of points on $C_{p}$ whose associated matrix $B_{k}(\mathbf{P})$ given by (1) has rank $2$ with the first $2$ columns of $B_{k}(\mathbf{P})$ being linearly independent and $m_{k,1}$ denote the number of element-wise distinct $k$-tuples where $B_{k}(\mathbf{P})$ has rank $1$.
Proposition 4.
For $k\geq 1$, we have
$\sum_{l_{1}\neq l_{2}\neq\cdots\neq l_{k}}E[Y_{l_{1}}Y_{l_{2}}\cdots Y_{l_{k}}]=m_{k,2}/p^{2}+m_{k,1}/p.$ (14)
Proof.
For any element-wise distinct tuple $l_{1},l_{2},\cdots,l_{k}$, it is easy to see that $P(Y_{l_{1}}=Y_{l_{2}}=\cdots=Y_{l_{k}}=1)=n/p^{2}$ where $n$ is the number of solutions $(i,j)$ of the equation
$B_{k}(\mathbf{P})\begin{pmatrix}i\\ j\\ 1\end{pmatrix}\equiv\begin{pmatrix}0\\ 0\\ \vdots\\ 0\end{pmatrix}\pmod{p}.$ (15)
In case the rank of $B_{k}(\mathbf{P})$ is $2$ and the first $2$ columns are linearly independent, there is a unique solution of (15) and hence $n=1$. In case the rank of $B_{k}(\mathbf{P})$ is $1$, there are $n=p$ distinct solutions of (15). In the other cases, either rank of $B_{k}(\mathbf{P})$ is $3$ or the rank of $B_{k}(\mathbf{P})$ is $2$ and the first $2$ columns are linearly dependent, it can be seen that there are no solutions of (15).
Since $E[Y_{l_{1}}Y_{l_{2}}\cdots Y_{l_{k}}]=P(Y_{l_{1}}=Y_{l_{2}}=\cdots=Y_{l_{k}}=1)$, equation (14) follows. ∎
Corollary 2.
For $k\geq 1$, and assuming that Conjecture 1 holds, we have
$\lim_{p\rightarrow\infty}\sum_{l_{1}\neq l_{2}\neq\cdots\neq l_{k}}E[Y_{l_{1}}Y_{l_{2}}\cdots Y_{l_{k}}]=1$
Proof.
Recall that in Section 1, we had stated that $m_{1,2}=0$ trivially and $m_{1,1}=m$. Hence by the Hasse-Weil bounds [10], we get that $m_{1,2}/p^{2}\rightarrow 0$ and $m_{1,1}/p\rightarrow 1$ as $p\rightarrow\infty$. Using this in equation (14), we get the required result for $k=1$.
Also for $k\geq 2$, we note that $m_{k,1}$ is bounded by a constant depending on $f$. This is because, for $k=2$, we can use Bezout’s theorem to bound $m_{k,2}$ by the number of solutions of the following system of $4$ equations in $4$ unknowns
$f(x,y),f(r,s)$ $\equiv 0\pmod{p};$
$f_{x}(x,y)f_{y}(r,s)-f_{x}(r,s)f_{y}(x,y)$ $\equiv 0\pmod{p};$
$f_{x}(x,y)f(r,s)-f_{x}(r,s)f(x,y)$ $\equiv 0\pmod{p^{2}}.$
For the case $k>2$, there are more equations than there are unknowns and in general there are no solutions. In either case, $m_{k,1}/p\rightarrow 0$ as $p\rightarrow\infty$. Using this and Conjecture 1 in equation (14), we get the required result for $k\geq 2$. ∎
4 Uniform distribution results
At the outset, we describe the version of uniform distribution that Conjecture 1 provably is a consequence of. First we give the following definition of discrepancy.
$\mathcal{D}_{p}(\mathcal{F}_{p})=\sup_{E\subseteq\mathbb{F}_{p}^{3}}\left|\frac{|\mathcal{F}_{p}\cap E|}{|E|}\frac{p^{3}}{m}-1\right|.$ (16)
Using this, we formulate the following conjecture on uniform distribution of $\mathcal{F}_{p}$.
Conjecture 2.
For $p\rightarrow\infty$, we have $\mathcal{D}_{p}(\mathcal{F}_{p})\rightarrow 0$.
We now give the proof of Conjecture 1 assuming this version of uniform distribution.
Proposition 5.
Conjecture 2 implies Conjecture 1.
Proof.
We start by noting that Conjecture 2 can be applied to any independent choice of $k$ vectors from $\mathcal{F}_{p}$. More precisely, Conjecture 2 implies
$\sup_{E\subseteq(\mathbb{F}_{p}^{3})^{k}}\left|\frac{|\mathcal{F}_{p}^{k^{\prime}}\cap E|}{|E|}\frac{p^{3k}}{m(m-1)\ldots(m-k+1)}-1\right|\rightarrow 0,$ (17)
as $p\rightarrow\infty$ where $\mathcal{F}_{p}^{k^{\prime}}$ deontes the element-wise distinct $k$-tuples of elements of $\mathcal{F}_{p}$.
Applying (17) to $E=\{(v_{1},v_{2},\ldots,v_{k})\in(\mathbb{F}_{p}^{3})^{k}|\ \textit{associatedmatrix has rank }2$ and first 2 columns linearly independent}}\}} gives
$\left|\frac{m_{k,2}}{|E|}\frac{p^{3k}}{m(m-1)\cdots(m-k+1)}-1\right|\rightarrow 0.$
Now $|E|=p^{3}(p^{3}-p)(p^{2}-2)(p^{2}-3)\cdots(p^{2}-k+1)$ and $m/p\rightarrow 1$ as $p\rightarrow\infty$ by the Hasse-Weil bounds [10]. Hence, we get (as $p\rightarrow\infty$)
$\left|\frac{m_{k,2}}{p^{2}}-1\right|\rightarrow 0,$ (18)
which is the same as (2). ∎
For comparison with Conjecture 2, we recall the standard version of discrepancy in the $3$-dimensional unit cube of the sequence of $\mathcal{F}_{p}$ scaled down by a factor of $p$ in each component, which we denote $\Delta_{p}(\mathcal{F}_{p})$.
$\Delta_{p}(\mathcal{F}_{p})=\sup_{E\subseteq\mathbb{F}_{p}^{3}}\left|\frac{|\mathcal{F}_{p}\cap E|}{m}-\frac{|E|}{p^{3}}\right|.$ (19)
We now state the standard version of uniform distribution of $\mathcal{F}_{p}$ as a conjecture.
Conjecture 3.
For $p\rightarrow\infty$, we have $\Delta_{p}(\mathcal{F}_{p})\rightarrow 0$.
Clearly, we have Conjecture 2 implies Conjecture 3, as $\Delta_{p}(\mathcal{F}_{p})\leq\mathcal{D}_{p}(\mathcal{F}_{p})$. It is also easy to see that Conjecture 3 gives $m_{2,2}/p^{2}\rightarrow 1$ as $p\rightarrow\infty$. However it is not yet clear whether it suffices to prove Conjecture 1 in its entirety.
On the other hand, we note the connection of Conjecture 3 with exponential sums by recalling the Erdös-Turan-Koksma inequality (Theorem 1.21, [5]).
Proposition 6.
[5] There exists a constant $C>0$ such that, for any positive integer $L$, the following bound holds
$\Delta_{p}(\mathcal{F}_{p})\leq C\left(\frac{1}{L}+\frac{1}{m}\sum_{0<|\mathbf{a}|\leq L}\frac{1}{r(\mathbf{a})}\left|\sum_{\mathbf{v}\in\mathcal{F}_{p}}e_{p}(\mathbf{a}\cdot\mathbf{v})\right|\right),$ (20)
where the sum runs over triples $\mathbf{a}\in\mathbb{Z}^{3}$ which satisfy the given condition with $|\mathbf{a}|=\max_{i=1}^{3}a_{i}$, $r(\mathbf{a})=\prod_{i=1}^{3}\max(|a_{i}|,1)$ and $e_{p}(x)=\exp(2\pi ix/p)$.
We finally state the following conjecture on exponential sums which appears to be obtainable along the same lines as Bombieri [1].
Conjecture 4.
There exists a constant $D>0$ such that, for any prime $p$ and any $k,l\in\{0,1,\dots,p-1\}$, the following bound holds
$\sum_{(x,y)\in C_{p}}e_{p^{2}}\left(f(x+kp,y+lp)\right)\leq Dp^{\frac{1}{2}}.$
It is easy to see by using Proposition 6 that Conjecture 4 implies that $\Delta_{p}(\mathcal{F}_{p})\ll p^{-\frac{1}{2}}$ which definitely implies Conjecture 3. However, it seems likely that Conjecture 2 is not a consequence of this estimate.
Numerical computations support the uniform distribution conjectures, but the speed of convergence appears to follow Conjecture 4. More precisely, we compute lower bounds for $\Delta_{p}(\mathcal{F}_{p}),\mathcal{D}_{p}(\mathcal{F}_{p})$ in Table 1 which are comparable with the order of decay predicted by Conjecture 4. These are obtained by restricting $E$ to cubes of side $p/3$ in (16),(19).
$p$ Lower bound for $\Delta_{p}$ Lower bound for $\mathcal{D}_{p}$ $p^{-1/2}$
$1009$ $0.01088009538$ $0.29376257545$ $0.03148142750$
$2003$ $0.01131027426$ $0.30537740503$ $0.02234392810$
$3001$ $0.00930693824$ $0.25128733264$ $0.01825437644$
$4003$ $0.00494879436$ $0.13361744798$ $0.01580546236$
$5443$ $0.00511927708$ $0.13822048125$ $0.01355441669$
$6007$ $0.00672228668$ $0.18150174042$ $0.01290242026$
$7013$ $0.00646586788$ $0.17457843276$ $0.01194120297$
$8011$ $0.00496738448$ $0.13411938098$ $0.01117266132$
$9007$ $0.00563262285$ $0.15208081695$ $0.01053682867$
$10009$ $0.00494777953$ $0.13359004739$ $0.00999550303$
Table 1: Lower bounds on discrepancy for various primes for the polynomial $f(x,y)=x^{3}+y^{2}+xy+1$.
5 Proof of Theorem 1
We start by noting that the random variable $X(k,l)$ defined in Section 1, representing the total number of points in $C_{p}$ with valuation greater than $1$ in the random $p\times p$ square, has the following expression
$X(k,l)=\sum_{(x,y)\in C_{p}}Y_{x,y}(k,l)=\sum_{i=1}^{m}Y_{i}(k,l).$ (21)
Recall that $m=|C_{p}|$.
By the properties of the random variables $Y_{x,y}$ discussed in Section 3, we have
$E[X]$ $=\sum_{i=1}^{m}E[Y_{i}]=m/p.$
In general for $k\geq 1$
$E[X^{k}]$ $=E\big{[}\big{(}\sum_{l=1}^{m}Y_{l}\big{)}^{k}\big{]}=E\big{[}\big{(}\sum_{l_{1},l_{2},\cdots,l_{k}=1}^{m}Y_{l_{1}}Y_{l_{2}}\ldots Y_{l_{k}}\big{)}\big{]}$
$=\sum_{i=1}^{k}S(k,i)\sum_{l_{1}\neq l_{2}\neq\cdots\neq l_{i}}E[Y_{l_{1}}Y_{l_{2}}\ldots Y_{l_{i}}],$ (22)
where $l_{i}\in\{1,2,...,m\}$ and $S(k,i)$ is the Stirling number of the second kind. This is because for each element-wise distinct tuple $(l_{1},l_{2},\cdots,l_{i})$ in the last sum, the contribution from each product, where only the variables $Y_{l_{1}},Y_{l_{2}},\ldots,Y_{l_{i}}$ appear, is the value $E[Y_{l_{1}}Y_{l_{2}}\ldots Y_{l_{i}}]$ Since the positions where these variables appear, form a partition of $\{1,2,\cdots,k\}$ into $i$ parts, we get $S(k,i)$ many such products for each $(l_{1},l_{2},\cdots,l_{i})$.
Thus, from (5) and Corollary 2 (assuming Conjecture 1 as in the theorem), we get the following limit for the $k^{\mathrm{th}}$ moment $E[X^{k}]$,
$\lim_{p\to\infty}E[X^{k}]=\sum_{i=1}^{k}S(k,i),$
which is indeed the $k^{\mathrm{th}}$ moment of the Poisson distribution with parameter $\lambda=1$ (see [8]). In other words, the moment generating function of $X$, $M_{X}(t)$, converges to
$\lim_{p\to\infty}M_{X}(t)=\exp(e^{t}-1).$
Hence, $X$ converges in distribution to a Poisson random variable with parameter $\lambda=1$ as $p\rightarrow\infty$.